OpenStudy (anonymous):

Find the x and y intercepts and graph the equation plotting the intercepts. -2x+7y=-7

OpenStudy (anonymous):

convert the equation to slope intercept form: divide through by 7: 2x / 7 + y = -1 y = -2x/7 - 1 y intercept is when x = 0 so y intercept is -1 x intercept is when y = 0 2x/7 = -1 x = -3.5

OpenStudy (anonymous):